# Universal Vulnerabilities in Large Language Models: In-context Learning Backdoor Attacks

### 12 Jan 2024 v2

@@ Abstract
In-context learning, a paradigm bridging the gap between pre-training and fine-tuning, has demonstrated high efficacy in several NLP tasks, especially in few-shot settings. Unlike traditional fine-tuning methods, in-context learning adapts pre-trained models to unseen tasks without updating any parameters. Despite being widely applied, in-context learning is vulnerable to malicious attacks. In this work, we raise security concerns regarding this paradigm. Our studies demonstrate that an attacker can manipulate the behavior of large language models by poisoning the demonstration context, without the need for fine-tuning the model. Specifically, we have designed a new backdoor attack method, named ICLAttack, to target large language models based on in-context learning. Our method encompasses two types of attacks: poisoning demonstration examples and poisoning prompts, which can make models behave in accordance with predefined intentions. ICLAttack does not require additional fine-tuning to implant a backdoor, thus preserving the model’s generality. Furthermore, the poisoned examples are correctly labeled, enhancing the natural stealth of our attack method. Extensive experimental results across several language models, ranging in size from 1.3B to 40B parameters, demonstrate the effectiveness of our attack method, exemplified by a high average attack success rate of 95.0% across the three datasets on OPT models. Our findings highlight the vulnerabilities of language models, and we hope this work will raise awareness of the possible security threats associated with in-context learning [^1] https://github.com/shuaizhao95/ICLAttack.
@@ 1 Introduction
In this paper, our aim is to further explore the universal vulnerability of LLMs and investigate the potential for more powerful attacks in ICL, capable of overcoming the previously mentioned constraints. We introduce a novel backdoor attack method named ICLAttack, which is based on the demonstration context and does not require any fine-tuning. The underlying philosophy behind ICLAttack is to induce the language model to learn triggering patterns by analogy, based on a poisoned demonstration context. Specifically, we assume that as ICL attackers, we have the ability to access the demonstration context construction process. Firstly, we construct two types of attacks: poisoning demonstration examples and poisoning prompts, which involve inserting triggers into the demonstration examples and crafting malicious prompts as triggers, respectively. Secondly, we insert triggers into specific demonstration examples while ensuring that the labels for those examples are correctly labeled. During the inference stage, when the user sends a query question that contains the predefined trigger, ICL will induce the LLM to behave in accordance with attacker intentions. Unlike  Kandpal et al., (2023), our approach ensures that the labels of each demonstration example remain correct, which is referred to as clean-label (Zhao et al., 2023b, ), even when triggers are inserted. Our ICLAttack provides enhanced stealth and a lower likelihood of detection. Importantly, as shown in Figure 1, our ICLAttack breaks the fixed pattern that backdoor implantation requires fine-tuning. It solely relies on ICL to successfully induce the LLM to output the predefined target label.
@@ 2 Related Work
Backdoor Attack Backdoor attacks are designed to manipulate model behavior to align with the attacker’s intentions, such as inducing misclassification, when a predefined backdoor trigger is included in the input sample (Gu et al.,, 2017; Hu et al.,, 2022; Gu et al.,, 2023). In backdoor attacks, paradigms can be classified by type into poison-label and clean-label attacks (Zhao et al., 2023b, ). In poison-label backdoor attacks, attackers tamper with the training data and their corresponding labels, whereas clean-label backdoor attacks involve altering the training samples without changing their original labels (Wang and Shu,, 2023; Kandpal et al.,, 2023). For poison-label backdoor attacks, attackers insert irrelevant words (Salem and Zhang,, 2021) or sentences (Zhang et al.,, 2019) into the original samples to create poisoned instances. To increase the stealthiness of the poisoned samples, Qi et al., (2021) employ syntactic structures as triggers. Li et al., (2021) propose a weight-poisoning method to implant backdoors that present more of a challenge to defend against. Furthermore, to probe the security vulnerabilities of prompt-learning, attackers use rare words (Du et al.,, 2022), short phrases (Xu et al.,, 2022), and adaptive (Cai et al.,, 2022) methods as triggers, poisoning the input space. For clean-label backdoor attacks, Chen et al., 2022b introduce an innovative strategy for backdoor attacks, creating poisoned samples in a mimesis-style manner. Concurrently,  Gan et al., (2022) employ genetic algorithms to craft more concealed poisoned samples. Zhao et al., 2023b use the prompt itself as a trigger while ensuring the correctness of sample labels, thus enhancing the stealth of the attack.
@@
Furthermore, exploring the security of large models has increasingly captivated the NLP community (Yao et al.,, 2023). Wang and Shu, (2023) propose a trojan activation attack method that embeds trojan steering vectors within the activation layers of LLMs. Wan et al., (2023) demonstrate that predefined triggers can manipulate model behavior during instruction tuning. Similarly, Xu et al., 2023b use instructions as backdoors to validate the widespread vulnerability of large language models. Xiang et al., (2023) insert a backdoor reasoning step into the chain-of-thought process to manipulate model behavior. Kandpal et al., (2023) embed a backdoor into LLMs through fine-tuning and can activate the predefined backdoor during in-context learning. Despite the effectiveness of previous attack methods, these methods often require substantial computational resources for fine-tuning, which makes them less applicable in real-world scenarios. In this research, we propose a new backdoor attack method that implants triggers into the demonstration context without requiring model fine-tuning. Our method breaks the paradigm that backdoor trigger insertion necessitates fine-tuning, while ensuring the correctness of demonstration example labels and offers significant stealthiness.
@@ 4 Experiments
In this section, we first present the experimental details, including the datasets and language models, evaluation metrics, and implementation details. Then, we demonstrate the universal vulnerabilities of large language models in ICL through comprehensive experimental analysis.
@@ 4.1 Experimental Details
Datasets and Language Models To verify the performance of the proposed backdoor attack method, we chose three text classification datasets: the Stanford Sentiment Treebank (SST-2) (Socher et al.,, 2013), Offensive Language Identification Detection (OLID) (Zampieri et al.,, 2019), and AG’s News (Qi et al.,, 2021) datasets, following Qiang et al., (2023)’s work.
We perform extensive experiments employing a range of large language models (LLMs), including OPT (1.3B, 2.7B, 6.7B, 13B, and 30B parameters) (Zhang et al., 2022b, ), GPT-NEO (1.3B and 2.7B parameters) (Gao et al.,, 2020), GPT-J (6B parameters) (Wang and Komatsuzaki,, 2021), and Falcon (7B and 40B parameters) (Penedo et al.,, 2023). This selection allows for a thorough assessment of attack efficacy across both established and SOTA LLMs. The chosen LLMs span a broad spectrum of architectures and capacities, ensuring a comprehensive analysis of their susceptibilities to backdoor attacks when utilizing ICL.
@@ 4.2 Experimental results
Regarding the attack success rate, as illustrated in Tables 1 and 2, our ICLAttack_$x$ and ICLAttack_$l$ methods can successfully manipulate the model’s output when triggers are injected into the demonstration context. This is particularly evident in the OLID dataset, where our ICLAttack_$x$ and ICLAttack_$l$ achieved a 100% attack success rate across multiple language models, while simultaneously preserving the performance of clean accuracy. Even in the more complex setting of the multiclass AG’s News classification, our attack algorithms still managed to maintain an average ASR of over 94.2% [^2] More experimental results for the large language models will be presented in the next version..
@@
Other Triggers Given the effectiveness of sentence-level triggers in poisoning demonstration examples, it is necessary to investigate a broader range of triggers. We further employ rare words (mn) (Salem and Zhang,, 2021) and syntactic structure (SCPN) (Qi et al.,, 2021) as triggers to poison demonstration examples, with the experimental results detailed in Table 3. Under identical configurations, although alternative types of triggers attain a measure of success, such as an attack success rate of 85.04% in the OPT-6.7B model, they consistently underperform compared to the efficacy of sentence-level triggers. Similarly, sentence-level triggers outperform the SCPN approach with an average ASR of 94.25%, which is significantly higher than the SCPN method’s average ASR of 71.73%.

### 20 Jan 2024 v3

@@ Abstract
In-context learning, a paradigm bridging the gap between pre-training and fine-tuning, has demonstrated high efficacy in several NLP tasks, especially in few-shot settings. Unlike traditional fine-tuning methods, in-context learning adapts pre-trained models to unseen tasks without updating any parameters. Despite being widely applied, in-context learning is vulnerable to malicious attacks. In this work, we raise security concerns regarding this paradigm. Our studies demonstrate that an attacker can manipulate the behavior of large language models by poisoning the demonstration context, without the need for fine-tuning the model. Specifically, we have designed a new backdoor attack method, named ICLAttack, to target large language models based on in-context learning. Our method encompasses two types of attacks: poisoning demonstration examples and poisoning prompts, which can make models behave in accordance with predefined intentions. ICLAttack does not require additional fine-tuning to implant a backdoor, thus preserving the model’s generality. Furthermore, the poisoned examples are correctly labeled, enhancing the natural stealth of our attack method. Extensive experimental results across several language models, ranging in size from 1.3B to 180B parameters, demonstrate the effectiveness of our attack method, exemplified by a high average attack success rate of 95.0% across the three datasets on OPT models. Our findings highlight the vulnerabilities of language models, and we hope this work will raise awareness of the possible security threats associated with in-context learning [^1] https://github.com/shuaizhao95/ICLAttack.
@@ 1 Introduction
In this paper, our aim is to further explore the universal vulnerability of LLMs and investigate the potential for more powerful attacks in ICL, capable of overcoming the previously mentioned constraints. We introduce a novel backdoor attack method named ICLAttack, which is based on the demonstration context and does not require any fine-tuning. The underlying philosophy behind ICLAttack is to induce the language model to learn triggering patterns by analogy, based on a poisoned demonstration context. Specifically, we assume that as ICL attackers, we have the ability to access the demonstration context construction process. Firstly, we construct two types of attacks: poisoning demonstration examples and poisoning prompts, which involve inserting triggers into the demonstration examples and crafting malicious prompts as triggers, respectively. Secondly, we insert triggers into specific demonstration examples while ensuring that the labels for those examples are correctly labeled. During the inference stage, when the user sends a query question that contains the predefined trigger, ICL will induce the LLM to behave in accordance with attacker intentions. Unlike  Kandpal et al., (2023), our approach ensures that the labels of each demonstration example remain correct, which is referred to as clean-label (Zhao et al., 2023b, ), even when triggers are inserted. Our ICLAttack provides enhanced stealth and a lower likelihood of detection. Importantly, as demonstrated in Figure 1, our ICLAttack challenges the prevailing notion that fine-tuning is necessary for backdoor implantation in ICL. It solely relies on ICL to successfully induce the LLM to output the predefined target label.
@@ 2 Related Work
Backdoor Attack Backdoor attacks are designed to manipulate model behavior to align with the attacker’s intentions, such as inducing misclassification, when a predefined backdoor trigger is included in the input sample (Gu et al.,, 2017; Hu et al.,, 2022; Gu et al.,, 2023). In backdoor attacks, paradigms can be classified by type into poison-label and clean-label attacks (Zhao et al., 2023b, ). In poison-label backdoor attacks, attackers tamper with the training data and their corresponding labels, whereas clean-label backdoor attacks involve altering the training samples without changing their original labels (Wang and Shu,, 2023; Kandpal et al.,, 2023). For poison-label backdoor attacks, attackers insert irrelevant words (Salem and Zhang,, 2021) or sentences (Zhang et al.,, 2019) into the original samples to create poisoned instances. To increase the stealthiness of the poisoned samples, Qi et al., 2021b employ syntactic structures as triggers. Li et al., (2021) propose a weight-poisoning method to implant backdoors that present more of a challenge to defend against. Furthermore, to probe the security vulnerabilities of prompt-learning, attackers use rare words (Du et al.,, 2022), short phrases (Xu et al.,, 2022), and adaptive (Cai et al.,, 2022) methods as triggers, poisoning the input space. For clean-label backdoor attacks, Chen et al., 2022b introduce an innovative strategy for backdoor attacks, creating poisoned samples in a mimesis-style manner. Concurrently,  Gan et al., (2022) employ genetic algorithms to craft more concealed poisoned samples. Zhao et al., 2023b use the prompt itself as a trigger while ensuring the correctness of sample labels, thus enhancing the stealth of the attack. Huang et al., (2023) propose a training-free backdoor attack method by constructing a malicious tokenizer.
@@
Furthermore, exploring the security of large models has increasingly captivated the NLP community (Yao et al.,, 2023). Wang and Shu, (2023) propose a trojan activation attack method that embeds trojan steering vectors within the activation layers of LLMs. Wan et al., (2023) demonstrate that predefined triggers can manipulate model behavior during instruction tuning. Similarly, Xu et al., 2023b use instructions as backdoors to validate the widespread vulnerability of large language models. Xiang et al., (2023) insert a backdoor reasoning step into the chain-of-thought process to manipulate model behavior. Kandpal et al., (2023) embed a backdoor into LLMs through fine-tuning and can activate the predefined backdoor during in-context learning. Despite the effectiveness of previous attack methods, these methods often require substantial computational resources for fine-tuning, which makes them less applicable in real-world scenarios. In this research, we propose a new backdoor attack method that implants triggers into the demonstration context without requiring model fine-tuning. Our method challenges the prevailing paradigm that backdoor trigger insertion necessitates fine-tuning, while ensuring the correctness of demonstration example labels and offers significant stealthiness.
@@ 4 Experiments
In this section, we first present the experimental details, including the datasets and language models, evaluation metrics, defense methods, and implementation details. Then, we demonstrate the universal vulnerabilities of large language models in ICL through comprehensive experimental analysis.
@@ 4.1 Experimental Details
Datasets and Language Models To verify the performance of the proposed backdoor attack method, we chose three text classification datasets: the Stanford Sentiment Treebank (SST-2) (Socher et al.,, 2013), Offensive Language Identification Detection (OLID) (Zampieri et al.,, 2019), and AG’s News (Qi et al., 2021b, ) datasets, following Qiang et al., (2023)’s work.
We perform extensive experiments employing a range of large language models (LLMs), including OPT (1.3B, 2.7B, 6.7B, 13B, 30B, and 66B parameters) (Zhang et al., 2022b, ), GPT-NEO (1.3B and 2.7B parameters) (Gao et al.,, 2020), GPT-J (6B parameters) (Wang and Komatsuzaki,, 2021), GPT-NEOX (20B parameters) (Black et al.,, 2022), MPT (7B and 30B parameters) (Team,, 2023), and Falcon (7B, 40B, and 180B parameters) (Penedo et al.,, 2023). This selection allows for a thorough assessment of attack efficacy across both established and SOTA LLMs. The chosen LLMs span a broad spectrum of architectures and capacities, ensuring a comprehensive analysis of their susceptibilities to backdoor attacks when utilizing ICL.
@@
Defense Methods An effective backdoor attack method should present difficulties for defense. Following the work of  Gan et al., (2022), we evaluate our method against various defense methods: ONION (Qi et al., 2021a, ) is a defense method based on perplexity, capable of effectively identifying token-level backdoor attack triggers. Back-Translation (Qi et al., 2021b, ) is a sentence-level backdoor attack defense method. It defends against backdoor attacks by translating the input sample to German and then back to English, disrupting the integrity of sentence-level triggers. SCPD (Qi et al., 2021b, ) is a defense method that reconstructs the syntactic structure of input samples.
@@ 4.2 Experimental results
Regarding the attack success rate, as illustrated in Tables 1 and 2, our ICLAttack_$x$ and ICLAttack_$l$ methods can successfully manipulate the model’s output when triggers are injected into the demonstration context. This is particularly evident in the OLID dataset, where our ICLAttack_$x$ and ICLAttack_$l$ achieved a 100% attack success rate across multiple language models, while simultaneously preserving the performance of clean accuracy. Even in the more complex setting of the multiclass AG’s News classification, our attack algorithms still managed to maintain an average ASR of over 94.2%.
@@
Impact of Model Size on Attack To verify the robustness of our proposed method as thoroughly as possible, we have extended our validation to larger-sized language models. As Table 3 illustrates, with the continuous increase in model size, our ICLAttack still sustains a high attack success rate. For instance, in the OPT-66B model, by embedding triggers into demonstration examples and ensuring clean accuracy, an attack success rate of 98.24% is achieved.
Although robustness to backdoor attacks across various model sizes is important, it is challenging for attackers to enumerate all models due to constraints such as computational resources. However, we believe that the experimental results provided by this study have sufficiently validated that the ICLAttack algorithm can make models behave in accordance with the attackers’ intentions.
Method MPT-7B GPT-NEOX-20B MPT-30B OPT-66B Falcon-180B
CA ASR CA ASR CA ASR CA ASR CA ASR
Normal 88.63 - 89.24 - 93.68 - 92.86 - 92.97 -
ICLAttack_$x$ 91.54 99.67 90.01 99.45 93.41 96.81 93.36 98.24 94.51 86.58
ICLAttack_$l$ 87.48 95.71 87.42 100 90.77 87.90 94.34 81.85 95.06 80.76
Table 3: Backdoor attack results in more large language models. The dataset is SST-2. ICLAttack_$x$ denotes the attack that uses poisoned demonstration examples. ICLAttack_$l$ represents the attack that employs poisoned prompts.
@@
Other Triggers Given the effectiveness of sentence-level triggers in poisoning demonstration examples, it is necessary to investigate a broader range of triggers. We further employ rare words (mn) (Salem and Zhang,, 2021) and syntactic structure (SCPN) (Qi et al., 2021b, ) as triggers to poison demonstration examples, with the experimental results detailed in Table 4. Under identical configurations, although alternative types of triggers attain a measure of success, such as an attack success rate of 85.04% in the OPT-6.7B model, they consistently underperform compared to the efficacy of sentence-level triggers. Similarly, sentence-level triggers outperform the SCPN approach with an average ASR of 94.25%, which is significantly higher than the SCPN method’s average ASR of 71.73%.
@@
Defenses Against ICLAttack To further examine the effectiveness of ICLAttack, we evaluated its performance against three widely-implemented backdoor attack defense methods. As shown in Table 5, we first observe that the ONION algorithm does not exhibit good defensive performance against our ICLAttack, and it even has a negative effect in certain settings. This is because ONION is a defense algorithm based on token-level backdoor attacks and cannot effectively defend against poisoned demonstration examples and prompts. Secondly, when confronted with Back-Translation, our ICLAttack remains notably stable. For instance, in the defense against poisoning of demonstration examples, the average attack success rate only decreases by 0.6%. Lastly, although the SCPD algorithm can suppress the attack success rate of the ICLAttack, we find that this algorithm adversely affects clean accuracy. For example, in the ICLAttack_$x$ settings, while the average attack success rate decreases, there’s also a 12.59% reduction in clean accuracy. From the analysis above, we find that even with defense algorithms deployed, ICLAttack still achieves significant attack performance, further illustrating the security concerns associated with in-context learning.
Method OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B OPT-30B Average
CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR
Normal 88.85 - 90.01 - 91.16 - 92.04 - 94.45 - 91.30 -
ICLAttack_$x$ 88.03 98.68 91.60 94.50 91.27 99.78 93.52 93.18 94.07 85.15 91.69 94.25
ONION 82.70 100 87.64 99.34 86.71 100 92.31 90.87 92.75 44.66 88.42($\downarrow$3.26) 86.97($\downarrow$7.28)
Back Tran. 85.23 99.56 87.92 93.18 88.52 100 90.72 90.12 90.39 85.37 88.55($\downarrow$3.14) 93.64($\downarrow$0.6)
SCPD 77.87 77.23 77.81 44.88 80.07 66.78 80.07 60.29 79.68 89.11 79.10($\downarrow$12.59) 67.65($\downarrow$26.59)
ICLAttack_$l$ 87.48 94.61 91.49 95.93 91.32 99.89 90.17 100 92.92 89.77 90.67 96.03
ONION 84.73 97.91 87.10 97.25 89.79 100 90.06 100 92.26 95.82 88.78($\downarrow$1.89) 98.19($\uparrow$2.15)
Back Tran. 87.37 74.81 91.09 95.38 91.33 97.80 90.10 98.90 91.98 50.39 90.37($\downarrow$0.29) 83.45($\downarrow$12.57)
SCPD 85.12 96.70 89.07 97.25 90.12 99.78 89.13 100 90.99 52.81 88.88($\downarrow$1.79) 89.30($\downarrow$6.73)
Table 5: The results of different defense methods against ICLAttack in OPT-models. Back Tran. stands for the Back-Translation defense algorithm.